# Dual Conditional Diffusion Models for Sequential Recommendation

### v1

@@ 4.1. Sequential Recommendation Diffusion Framework
where $\phi$ represents the learnable parameters in $\text{Emb}(z)$. As for the reverse process, we define the predicted distribution of $z$ as:
@@
where $\text{cos}(\cdot)$ is the cosine similarity between $e_{0}$ and each item embedding of $\text{Emb}(\mathbb{Z})$. However, the transition distribution $p_{\phi}(z|e_{0})$ lacks a direct analytical formula.
After defining $p_{\phi}$ and $q_{\phi}$, we extend the ELBO in Equation 7 to include discrete item $z$ as:
(13) $\arg\max\limits_{\theta}\log p_{\theta}(z)$ $\approx\arg\min\limits_{\phi,\theta}\mathbb{E}_{q_{\phi}(e_{0:T}|z)}\left[\log\frac{q_{\phi}(e_{0:T}|z)}{p_{\phi,\theta}(z,e_{0:T})}\right],$
and the training objective in Equation 9 is rewritten as:
@@
There are three terms of objectives in $\mathcal{L}_{\text{elbo}}$: the prior matching term $\mathcal{L}_{T}$, the denoising term $\mathcal{L}_{t}$ and the domain transition term $\mathcal{L}_{z}$. The prior matching term $\mathcal{L}_{T}$ is similar to the same term in Equation 9, which requires the forward process to transit the $e_{0}$ into Gaussian noise $p_{\theta}(e_{T})$. Continuous diffusion models (Luo, 2022; Ho et al., 2020; Ho and Salimans, 2022a; Kong et al., 2021) omit this term since there are no associated trainable parameters. This term is essential in $\mathcal{L}_{\text{elbo}}$ due to the correlation among $z$, $e_{0}$ and $\phi$. This term is equivalent to minimizing $\|e_{0}\|^{2}$. As the second term $\mathcal{L}_{t}$ require $f_{\theta}(e_{t},t)$ and $e_{0}$ as close as possible, we transfer the objective $\|e_{0}\|^{2}$ to $\|\sum_{t=1}^{T}f_{\theta}(e_{t},t)\|^{2}$ for fully optimize the DNN models.
The embedding entropy term $\mathcal{L}_{z}$ can be regarded as a KL-divergence $D_{\text{KL}}(q_{\phi}(e_{0}|z)||p_{\phi}(z|e_{0}))$. Although $q_{\phi}(e_{0}|z)$ is a linear transition, the distribution of $p_{\phi}(z|e_{0})$ is unknown. Therefore, we replace $D_{\text{KL}}(q_{\phi}(e_{0}|z)||p_{\phi}(z|e_{0}))$ by $D_{\text{KL}}(q(z)||p_{\phi}(z))=D_{\text{KL}}(q(z)||p_{\phi}(z))$. The simplified objective of $\mathcal{L}_{\text{elbo}}$ is:
@@
(17) $\mathcal{L}_{T}$ $=\|f_{\theta}(e_{t},t)\|^{2},$
(18) $\mathcal{L}_{t}$ $=\|f_{\theta}(e_{t},t)-e_{0}\|^{2},$
(19) $\mathcal{L}_{z}$ $=D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(z))$
@@ 4.2. Dual Conditional Diffusion Transformer
In this section, we provide a detailed explanation of the Dual Conditional Diffusion Transformer (DCDT) in DCRec, which is a robust denoising transformer built within the framework in Section 4.1. Previous methods in diffusion-based SR have primarily employed either implicit or explicit conditioning mechanisms to incorporate user behavioral data into the recommendation process (Yang et al., 2023b; Li et al., 2024a; Xie et al., 2024). Almost implicit conditional approaches compress the user’s historical behavior sequence into a compact representation vector, which is then used to guide the reverse diffusion process (Lin et al., 2024b; Li et al., 2024a). While this method efficiently captures global user preferences, it may inadvertently oversimplify complex user behaviors, losing important sequential and temporal dynamics inherent in the user’s interactions. This compression can lead to a loss of valuable information about the order and context of past actions, which are crucial for accurate SRs. Explicit conditioning methods (Lin et al., 2024a), on the other hand, skip the feature compression step and directly introduce the entire sequence of historical item embeddings during the reverse generation process. This allows the model to access detailed step-by-step interactions and capture fine-grained sequential dependencies. However, explicit conditioning on the whole historical information, these methods can be more sensitive to noise and irrelevant actions within the user’s behavior sequence, which may hinder the recommendation quality (He and McAuley, 2016a, b).
To tackle these limitations, we propose the Dual Conditioned Diffusion Transformer (DCDT) as shown in Figure 3, a novel approach that integrates both implicit and explicit conditioning strategies. By combining the representation of the user’s global preferences with the direct utilization of detailed historical behaviors, DCDT  leverages the strengths of both methods. The implicit component captures global user interests and long-term behavioral patterns, providing a stable guide for the diffusion process. Simultaneously, the explicit component preserves sequential complexities and temporal dynamics by referencing specific past interactions during generation. DCDT  employs a dual conditional diffusion process that incorporates an implicit conditional historical forward process and an explicit conditional reverse process, leading to more accurate recommendations. As shown in Figure 2(b), the dual conditional trajectory (green line) in the reverse diffusion process merges these two guiding signals, preventing the denoising trajectory from being overly influenced by noise or irrelevant historical items. This effectively narrows the trajectory toward the correct target item, improving the precision of the recommendation.
4.2.1. Implicit Conditional Historical Forward Process:
Previous implicit conditional methods tend to encode the user’s historical behavior sequences into compact feature representations, which are then used to guide the reverse diffusion process (Yang et al., 2023b; Li et al., 2024a). However, compressing the historical behavior sequences into a single vector might lead to the loss of important sequential information due to the sparse nature of SRSs (He and McAuley, 2016b). In our approach, we avoid this compression step. Instead, we combine the uncompressed historical behavior sequences with the target item and add noise to this combined input before feeding it into the diffusion model. We believe that this allows the diffusion model to generate a more robust representation of the target item, as it can better capture detailed sequential patterns and contextual information from the user’s history. For convenience, for our implicit conditional diffusion models, we use $S_{0}$ to donate the concatenated sequence of the history item sequence $H_{0}$ and the target $e_{0}$, where $S_{0}=\text{Concat}(H_{0},e_{0})$. Thus, we continuously incorporate Gaussian noises into the whole item sequences $S_{0}$ rather than the target item embedding $e_{0}$ as follows:
(20) $q(S_{t}|S_{t}-1)$ $=\mathcal{N}(S_{t};\sqrt{\alpha_{t}}S_{t-1},(1-\alpha_{t})\mathbf{I})$
(21) $=\sqrt{\alpha_{t}}S_{0}+\sqrt{1-\alpha_{t}}S_{t-1}$
Based on the superposition of Gaussian distribution and Equation 21, we can calculate the transition from $S_{0}$ to $S_{t}$ as:
(22) $q(S_{t}|S_{0})=\sqrt{\bar{\alpha}_{t}}S_{0}+\sqrt{1-\bar{\alpha}_{t}}S_{T},\ S_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, and $S_{t}$ denotes the item embeddings in the forward step $t$. We further add explicit conditioning by adding $H_{0}$ with a scaling factor $\delta$ to $S_{T}$ and rewrite Equation 22 as:
(23) $q(S_{t}|S_{0})=\sqrt{\bar{\alpha}_{t}}S_{0}+\sqrt{1-\bar{\alpha}_{t}}(S_{T}+\delta H_{0}),$
Therefore, in the ELBO-based training objective, we can rewrite Equation 17 and Equation 18 as:
[…]
Compared to existing implicit diffusion models-based SR methods (Yang et al., 2023b), our method adds noising the entire user sequence rather than just the target item. This approach allows the following reverse process to explicitly leverage the user’s interaction history when generating the target item that aligns with the user’s interests, leading to more accurate and relevant recommendations.
Figure 3. Our DCDT  comprises explicit conditional layer normalization, multi-head self-attention, and an explicit cross-attention.
4.2.2. Explicit Conditional Reverse Process:
In the context of SR, the restoration process reverses the transition modeling by gradually eliminating the noise and recovering the original sequential patterns of user behavior, thereby restoring the data distribution underlying the user’s item preferences. Meanwhile, cross attention as an explicit conditional method has proved effective in classic conditional generation tasks (Rombach et al., 2022; Podell et al., 2024). Therefore, we further introduce the user’s item preferences $H_{0}$ as an explicit conditional signal to the denoising transformer $f_{\theta}$ in Equation 24 as:
(25) $\hat{S}_{0}=f_{\theta}(S_{t},t,H_{0})$
@@ 4.2.2. Explicit Conditional Reverse Process:
Firstly, we introduce the Conditional Layer Normalization, abbreviated as CondLN. In our SR setting, we hope to extend traditional layer normalization to our explicit conditional denoising process. Therefore, we replace the element-wise learnable parameters gain $g$ and bias $b$ of affine transformation in the original Layer Normalization (Lei Ba et al., 2016) by integrating user interaction sequence $S_{0}$ and diffusion step embedding $t$. The CondLN can be formulated as:
(26) $\bar{S}_{t}=g^{\prime}(t,H_{0})\times\frac{S_{t}-\mathbb{E}(S_{t})}{\sqrt{\text{Var}(S_{t})+\epsilon}}+b^{\prime}(t,H_{0})$
@@
where $\mathbb{E}(\cdot)$ and $\text{Var}(\cdot)$ denote the mean and standard variance of $S_{t}$, and $\epsilon$ denotes the small constant for numerical stability. $g^{\prime}(\cdot)$, $b^{\prime}(\cdot)$ are the MLP function conditioned by $H_{0}$ and $t$. We scale the normalized $\text{Var}(S_{t})$ and shift it with $\mathbb{E}(S_{t})$.
@@
In this way, CondLN explicitly affects the generation process with user interaction history $H_{0}$. This allows for more adaptive and context-aware normalization compared to standard layer normalization, which is crucial for capturing the intricate dependencies in SR tasks. Next, we input this sequence embedding $\bar{S}_{t}$ into a self-attention module (SA):
@@
(27) $\mathbf{\bar{S}_{t}}=\text{SA}(\bar{S}_{t})=\text{Softmax}\left(\frac{(\bar{S}_{t}\mathbf{W}^{Q})(\bar{S}_{t}\mathbf{W}^{K})^{\top}}{\sqrt{d}}\right)(\bar{S}_{t}\mathbf{W}^{V})\in\mathbb{R}^{(n-1)\times d}$
where $\mathbf{W}^{Q}$,$\mathbf{W}^{K}$,$\mathbf{W}^{V}$ are learnable weight parameters, $d$ denotes the dimension of embedding. After obtaining the self-attention output, then we leverage the Cross-attention layer (CA):
(28) $\bar{S}_{t}=\text{CA}(\bar{S}_{t},\mathbf{H_{0}})=\text{Softmax}\left(\frac{(\bar{S}_{t}\mathbf{W}^{Q})(\mathbf{H_{0}}\mathbf{W}^{K})^{\top}}{\sqrt{d}}\right)(\mathbf{H_{0}}\mathbf{W}^{V})$
@@
where $\mathbf{W}^{Q}$, $\mathbf{W}^{K}$, $\mathbf{W}^{V}$ are learnable parameters. Due to the sparse interaction in SR (He and McAuley, 2016b), we adapt $\bar{S}_{t}$ as query and the $H_{0}$ (pure historical embeddings) as key, value. $H_{0}$ can provide more rich user interaction information than ambiguous $H_{t}$. Hence, we select a cross-attention architecture as the denoising decoder instantiation capable of taking the entire sequence representation and corresponding step indicator $t$ as input. Hence, we adopt a cross-attention architecture for the denoising decoder, which effectively processes the entire sequence representation along with the corresponding diffusion step indicator $t$ as input. Then we endow the traditional layer normalization and feed-forward network to obtain the final user sequence embedding $\hat{S}_{0}$:
(29) $\bar{S}_{t}=\text{LayerNorm}(\bar{S}_{t}+\text{Dropout}(\text{sublayer}(\bar{S}_{t})))$
(30) $\hat{S}_{0}=\text{FFN}(\bar{S}_{t})=\text{RELU}(\bar{S}_{t}\times\mathbf{W}+\mathbf{b})$
@@
After obtaining the $\hat{S}_{0}$, we extract the last item embedding $e_{0}$ of $\hat{S}_{0}$ as the generative target item embedding. Then we leverage the variant of Equation 12, to transit continuous $\hat{S}_{0}$ into the discrete item space $\mathbb{Z}$ as $z$, taking the item index $z$ corresponding to the maximal as the final recommendation list to the specific user. Finally, the optimization object of the dual conditional diffusion process can be reformulated as:
@@
(32) $\mathcal{L}_{T}$ $=\|f_{\theta}(e_{t},t)\|^{2},$
(33) $\mathcal{L}_{t}$ $=\|f_{\theta}(e_{t},t)-e_{0}\|^{2},$
(34) $\mathcal{L}_{z}$ $=D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(z))$
where $\lambda$ donates the balance factor mediating the focus between $\mathcal{L}_{t}$ and $\mathcal{L}_{z}$. Algorithm 1 in Appendix A provides a detailed explanation of DCRec’s training process.
4.3. Model Inference and Inference Acceleration
In the Section 4.1 and Section 4.2, we have introduced the optimization process in DCRec  and the training of DCDT. The denoising model DCDT  is trained to provide predicted $\hat{S}_{0}$ to approximate $S_{0}$. Similar to Equation 2, the inference Markov transition of DCRec  can be formulated as:
(35) $p_{\theta}(S_{t-1}|S_{t})=\mathcal{N}(S_{t-1};\mu_{\theta}(\hat{S}_{0}(S_{t}),t),\Sigma_{\theta}(\hat{S}_{0}(S_{t}),t))$
This can be described as a repeated approximation-and-adjust process since each step aims to predict $\hat{S}_{0}(S_{t})$ and provide $S_{t-1}$, which is less noise than $S_{t}$, for the next step. The time cost of inference process linearly increases associated with the increase of the total number of steps. A natural idea for acceleration is that, if the approximation at the early is accurate enough with no need for adjustment, DCRec  can directly skip several intermediate steps. Here, we define the number of skipping steps as k. We rewrite Equation 35 as:
(36) $p_{\theta}(S_{t-k}|S_{t})=\mathcal{N}(S_{t-k};\mu_{\theta}(\hat{S}_{0}(S_{t}),t),\Sigma_{\theta}(\hat{S}_{0}(S_{t}),t))$
For our DCRec, we observe that DCDT  is robust enough to fulfill Equation 36 at an early stage, which significantly reduces time overhead. We will discuss the details in Section 5.4. For the pseudo-code of the inference, please refer to Algorithm. 2 in the Appendix.
5. Experiment

### v2

@@ 2.2.2. Diffusion Models for Sequential Recommendation
where $\phi$ represents the learnable parameters in $\text{Emb}(z)$. In this paper, the corresponding reverse process, which maps vectors in the embedding space back to items is defined as follows:
@@
where $\text{cos}(\cdot)$ is the cosine similarity between $e_{0}$ and each item embedding of $\text{Emb}(\mathbb{Z})$. After defining $p_{\phi}$ and $q_{\phi}$, the training objective in Equation 3 can be extended to including $z$ as:
@@
There are three terms of objectives: the prior matching term $\mathcal{L}_{T}$, the denoising term $\mathcal{L}_{t}$ and the domain transition term $\mathcal{L}_{z}$. Different diffusion-based works may have different approaches to optimize $\mathcal{L}_{\text{elbo}}$, while in this paper, our optimization is based on (Li et al., 2022; Gong et al., 2023).
Specifically, The $\mathcal{L}_{T}$ is similar to the same term in Equation 3, which requires the forward process to transit the $e_{0}$ into Gaussian noise $p_{\theta}(e_{T})$. While vanilla diffusion models (Luo, 2022; Ho et al., 2020; Ho and Salimans, 2022a; Kong et al., 2021) omit this term since there are no associated trainable parameters, this term is essential in Equation 7, since $e_{0}$ are calculated from $z$ and $\phi$. Following (Li et al., 2022), this term can be simplified as $\|e_{0}\|^{2}$. We further transfer the objective $\|e_{0}\|^{2}$ to $\sum_{t=1}^{T}\|f_{\theta}(e_{t},t)\|^{2}$ since the second term $\mathcal{L}_{t}$ requires $f_{\theta}(e_{t},t)$ and $e_{0}$ as close as possible. As for $\mathcal{L}_{z}$, it can be regarded as a KL-divergence $D_{\text{KL}}(q_{\phi}(e_{0}|z)||p_{\phi}(z|e_{0}))$. However, since $p_{\phi}(z|e_{0})$ is a argmax function that lacks a direct analytical formula, we approximate $D_{\text{KL}}(q_{\phi}(e_{0}|z)||p_{\phi}(z|e_{0}))$ by $D_{\text{KL}}(q(z)||p_{\phi}(z))$, where $q(z)$ is the ground truth distribution and $p_{\phi}(z)$ is the generated distribution of item. Thus, the simplified objective of $\mathcal{L}_{\text{elbo}}$ is:
@@
(10) $\mathcal{L}_{T}$ $=\textstyle\sum\nolimits_{t=1}^{T}\|f_{\theta}(e_{t},t)\|^{2},$
(11) $\mathcal{L}_{t}$ $=\textstyle\sum\nolimits_{t=1}^{T}\|f_{\theta}(e_{t},t)-e_{0}\|^{2},$
(12) $\mathcal{L}_{z}$ $=D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(z)),$
3. Dual Conditional Diffusion Model
In this section, we present a comprehensive introduction of our Dual Conditional Diffusion Model (DCRec) as shown in Figure 2. First, we introduce the theoretical derivation of the implicit conditional forward process and the explicit conditional reverse process. Then, to bridge the gap between diffusion theory and a practical model, we introduce the Dual Conditional Diffusion Transformer (DCDT), our denoising network $f_{\theta}$ for diffusion SR is shown in Figure 3. Finally, we explore efficient inference for DCRec.
3.1. Implicit Conditional Forward Process
Implicit conditioning provides indirect guidance for the denoising model by integrating the user’s historical information into the noisy input data. Due to the input being calculated by Equation 2, this conditioning effects the model in the diffusion forward process. Previous implicit conditional methods tend to encode the user’s historical behavior sequences into compact feature representations, which are then used to guide the reverse diffusion process (Yang et al., 2023b; Li et al., 2024a). However, compressing the historical behavior sequences into a single vector might lead to the loss of important sequential information due to the sparse nature of SRSs (He and McAuley, 2016b).
In our approach, we avoid this compression by concatenating the uncompressed historical behavior sequences with the target item and adding noise to the sequence. We believe that this allows the diffusion model to generate a more robust representation of the target item, as it can better capture detailed sequential patterns and contextual information from the user’s history. Here, we use $S_{0}$ to denote the concatenated sequence of the history item sequence $H_{0}$ and the target $e_{0}$, where $S_{0}=\text{Concat}(H_{0},e_{0})$. The input of the denoising network $f_{\theta}$ is expanded from $e_{0}$ to $S_{0}$. Thus, we rewrite Equation 2 as follows:
(13) $q(S_{t}|S_{t-1})$ $=\mathcal{N}(S_{t};\sqrt{\alpha_{t}}S_{t-1},(1-\alpha_{t})\mathbf{I}),$
Meanwhile, the denoising DNN $f_{\theta}$ in the reverse process is extended to support sequential input as follows:
(14) $\bar{S}_{0}=f_{\theta}(S_{t},t),$
where $\bar{S}_{0}$ is the recovered target item.
Figure 3. The design of DCDT. The black box refers to the main architecture, the red box refers to the details within the transformer block, and the yellow box refers to the details of the CondLN module.
3.2. Explicit Conditional Reverse Process
Explicit conditioning provides direct guidance for the denoising model by influencing the reverse denoising process in Equation 1, thereby recovering the target item underlying the user’s preferences. Although this conditioning has proven promising in image synthesis tasks (Rombach et al., 2022; Podell et al., 2024), most recent diffusion-based SR methods (Yang et al., 2023b; Li et al., 2024b) primarily focus on implicit conditioning, without fully exploring the potential of explicit conditioning. Unlike implicit conditioning which may lose part of the sequential information in the user’s behavior sequence, explicit conditioning directly introduces the entire sequence of historical item embeddings.
After expanding the concatenated input $S_{t}$ in Section 3.1, we rewrite standard reverse transition Equation 1 with item preference $H_{0}$ as an explicit conditional signal for the denoising process:
(15) $p_{\theta}(S_{t-1}|S_{t},H_{0})=\mathcal{N}(S_{t-1};\mu_{\theta}(S_{t},t,H_{0}),\Sigma_{\theta}(S_{t},t,H_{0})),$
Combining implicit and explicit conditioning, the denoising DNN $f_{\theta}$ in Equation 14 can be further rewritten as:
(16) $\bar{S}_{0}=f_{\theta}(S_{t},t,H_{0}),$
After deriving the forward and reverse process as Equations 13 and 15, implementing a practical model remains challenging. In the next section, we introduce a novel network architecture designed to effectively integrate both types of conditioning.
@@ 3.3. Dual Conditional Diffusion Transformer
We introduce the Dual Conditioned Diffusion Transformer (DCDT) in DCRec, a robust denoising network incorporating both implicit and explicit conditioning.
@@
Previous implicit conditional approaches (Yang et al., 2023b; Li et al., 2024a; Xie et al., 2024) compress the user’s historical behavior sequence into a compact representation vector, which is then used to guide the reverse diffusion process (Lin et al., 2024b; Li et al., 2024a). While this method efficiently captures global user preferences, it may inadvertently oversimplify complex user behaviors, losing important sequential and temporal dynamics inherent in the user’s interactions. This compression can lead to a loss of valuable information about the order and context of past actions, which are crucial for accurate SRs. Explicit conditioning methods, on the other hand, skip the feature compression step and directly introduce the entire sequence of historical item embeddings during the reverse generation process. This allows the model to access detailed step-by-step interactions and capture fine-grained sequential dependencies. However, explicit conditioning on the whole historical information can be more sensitive to noise and irrelevant actions within the user’s behavior sequence, which may hinder recommendation quality (He and McAuley, 2016a, b).
To address these limitations, we jointly leverage both implicit and explicit conditioning. We propose DCDT. Figure 3 shows an overview of the complete DCDTarchitecture. In the following, we describe the architecture design and the loss function of DCDT.
3.3.1. Input Concatenation
The main input to DCDT  is a sequential noisy embedding $S_{t}$, concatenated with noisy user’s history sequence $S_{t}$ and noisy target item $e_{t}$. This concatenation provides implicit guidance for recover the pure target item. Following the concatenation, we apply a standard positional embedding as other transformer-based SRs (Kang and McAuley, 2018; Hidasi and Karatzoglou, 2018; Yuan et al., 2019; Sun et al., 2019; Zhou et al., 2020a).
3.3.2. Conditional Layer Normalization (CondLN):
@@
In our SR setting, we hope to extend traditional layer normalization to our explicit conditional denoising process. Therefore, we replace the element-wise learnable parameters gain $g$ and bias $b$ of affine transformation in the original Layer Normalization (Lei Ba et al., 2016) by integrating user interaction sequence $S_{0}$ and diffusion step embedding $t$. CondLN can be formulated as:
(17) $\bar{S}_{t}=g^{\prime}(t,H_{0})\times\frac{S_{t}-\mathbb{E}(S_{t})}{\sqrt{\text{Var}(S_{t})+\epsilon}}+b^{\prime}(t,H_{0}),$
@@
where $\mathbb{E}(\cdot)$ and $\text{Var}(\cdot)$ denote the mean and variance of $S_{t}$, and $\epsilon$ denotes a small constant for numerical stability. $g^{\prime}(\cdot)$, $b^{\prime}(\cdot)$ are MLP functions conditioned by $H_{0}$ and $t$. We scale the normalized $\text{Var}(S_{t})$ and shift it with $\mathbb{E}(S_{t})$.
@@
In this way, CondLN explicitly affects the generation process with the user interactions $H_{0}$. This allows for more adaptive and context-aware normalization compared to standard layer normalization, which is crucial for capturing the intricate dependencies.
3.3.3. Implicit Self-Attention:
Following the CondLN, we input the sequence embedding $\bar{S}_{t}$ into a self-attention module (SA):
@@ 3.3.3. Implicit Self-Attention:
(18) $\bar{S}_{t}=\text{SA}(\bar{S}_{t})=\text{Softmax}\left(\frac{(\bar{S}_{t}\mathbf{W}^{Q})(\bar{S}_{t}\mathbf{W}^{K})^{\top}}{\sqrt{d}}\right)(\bar{S}_{t}\mathbf{W}^{V})\in\mathbb{R}^{(n-1)\times d},$
where $\mathbf{W}^{Q}$,$\mathbf{W}^{K}$,$\mathbf{W}^{V}$ are learnable weight parameters, $d$ denotes the dimension of embedding. The self-attention module captures sequential features from the noisy input $\bar{S}_{t}$ and extracts the high-level representations implicit in the user preference. Since there is no extra conditioning input, this module is indirectly guided by the input sequence.
3.3.4. Explicit Cross-Attention:
After obtaining the self-attention output and applying layer normalization, we utilize the cross-attention module (CA) with the pure historical embedding $H_{0}$:
(19) $\bar{S}_{t}=\text{CA}(\bar{S}_{t},\mathbf{H_{0}})=\text{Softmax}\left(\frac{(\bar{S}_{t}\mathbf{W}^{Q})(\mathbf{H_{0}}\mathbf{W}^{K})^{\top}}{\sqrt{d}}\right)(\mathbf{H_{0}}\mathbf{W}^{V}),$
@@ 3.3.4. Explicit Cross-Attention:
where $\mathbf{W}^{Q}$, $\mathbf{W}^{K}$, $\mathbf{W}^{V}$ are learnable parameters. Due to the sparse interactions in SR (He and McAuley, 2016b), we adapt $\bar{S}_{t}$ as a query and $H_{0}$ as both key and value. $H_{0}$ can explicitly provide more rich user interaction information than ambiguous $H_{t}$. Hence, we adopt a cross-attention architecture for the denoising decoder, which effectively processes the entire sequence representation along with the corresponding diffusion step indicator $t$ as input.
3.3.5. Traditional Feed-forward:
Then we employ another layer normalization and the traditional feed-forward module to obtain the final sequence embedding of the DCDT  block:
(20) $\bar{S}_{t}=\text{FFN}(\bar{S}_{t})=\text{RELU}(\bar{S}_{t}\times\mathbf{W}+\mathbf{b}),$
where $\mathbf{W}$ and $\mathbf{b}$ are learnable parameters.
3.3.6. Network Output and Loss Function:
@@
After the repeated DCDT  blocks and the final CondLN module, we obtain the predicted target $\bar{S}_{0}=\text{Concat}(\bar{H}_{0},\bar{e}_{0})$. In training, we extract the last item embedding $\bar{e}_{0}$ of $\bar{S}_{0}$ as the generative target item embedding. Then we leverage the variant of Equation 6, to transit continuous $\bar{S}_{0}$ into the discrete item space $\mathbb{Z}$ as $\bar{z}$, taking the item index $\bar{z}$ corresponding to the maximal probability as the final recommendation list to the specific user. Finally, the standard optimization object Equation 9 can be reformulated for our dual conditional diffusion as:
@@
(22) $\mathcal{L}_{T}$ $=\|\bar{S}_{0}\|^{2},$
(23) $\mathcal{L}_{t}$ $=\|\bar{e}_{0}-e_{0}\|^{2},$
(24) $\mathcal{L}_{z}$ $=D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(\bar{z})),$
where $\lambda$ donates the balance factor between $\mathcal{L}_{t}$ and $\mathcal{L}_{z}$. Algorithm 1 provides details of the training process in DCRec.
Algorithm 1 DCRec  Training
2:repeat
3:     Sample ${H_{0}}$ from $\{H_{0}\}$ and corresponding $z$ from $\mathbb{Z}$
4:     $S_{0}$leftarrow Concatenate $H_{0}$ and $z$
5:     Sample $t\sim\mathcal{U}(1,T)$
6:     $S_{t}$leftarrow Adding noise to $S_{0}$ according to Equation 2
7:     Obtain predicted $\bar{S}_{0}\leftarrow f_{\theta}(S_{t},t,H_{0})$
8:     Extract predicted $\bar{e}_{0}$ from $\bar{S}_{0}$
9:     Obtain predicted $\bar{z}\leftarrow p_{\phi}(\bar{z}|\bar{e}_{0})$
10:     $\mathcal{L}_{T}$leftarrow $\|\bar{S}_{0}\|^{2}$
11:     $\mathcal{L}_{t}$leftarrow $\|\bar{e}_{0}-e_{0}\|^{2}$
12:     $\mathcal{L}_{z}$leftarrow $D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(\bar{z}))$
14:until converged
15:optimized $\theta$, $\phi$.
3.4. Model Inference and Acceleration
In the Section 3.3, DCDT  is trained to provide predicted $\bar{S}_{0}$ to approximate $S_{0}$. During model inference, DCDT  iteratively generates $S_{0}$ according to the total steps $T$. $\bar{S}_{0}$ will be transfer to $S_{t-1}$ for the next time step, formulated as $S_{t-1}=\bar{S}_{0}(S_{t})$. Based on Equation 15, the inference Markov transition of DCRec  can be reformulated as:
(25) $p_{\theta}(S_{t-1}|S_{t})=\mathcal{N}(S_{t-1};\mu_{\theta}(\bar{S}_{0}(S_{t}),t),\Sigma_{\theta}(\bar{S}_{0}(S_{t}),t))$
This can be described as a repeated approximation-and-adjustment process since each step aims to generate $\bar{S}_{0}$ base on the output from the last step. $\bar{S}_{0}$ becomes increasingly precise as inference progresses. The time cost of the inference process increases linearly in the total number of steps. A natural idea for acceleration is that if the approximation at the early stage is accurate enough with no need for adjustment, DCRec  can directly skip several intermediate steps according to the high similarity of $\bar{S}_{0}$. Here, we define the number of skipping steps as k. We rewrite Equation 25 as:
(26) $p_{\theta}(S_{t-k}|S_{t})=\mathcal{N}(S_{t-k};\mu_{\theta}(\bar{S}_{0}(S_{t}),t),\Sigma_{\theta}(\bar{S}_{0}(S_{t}),t)).$
For our DCRec, we observe that DCDT  is robust enough to fulfill Equation 26 at an early stage, which significantly reduces time overhead. We will discuss the details in Section 4.5. Algorithm 2 shows the pseudo-code of the inference procedure.
Algorithm 2 DCRec  Inference
2:Uniformly sample $T^{\prime}$ steps from $1$ to $T$
3:Sample $e_{T}~{}\sim\mathcal{N}(\bm{0},\bm{I})$
4:Sample $H_{T}~{}\sim\mathcal{N}(\bm{0},\bm{I})$
5:Implicit conditioning $H_{T}\leftarrow H_{T}+\delta H_{0}$
6:$S_{T}\leftarrow$ Concatenate $H_{T}$ and $e_{T}$
7:for $t\leftarrow T^{\prime}\dots 1$  do
8:     Obtain predicted ${S}_{0}\leftarrow f_{\theta}(S_{t},t,H_{0})$
9:     Transit to $S_{t-1}\leftarrow q(S_{t-1}|{S}_{0})$
10:Extract $e_{0}$ from $S_{0}$
11:obtain $z\leftarrow p_{\phi}(z|e_{0})$
12:$z$
4. Experiments